OpenStudy (anonymous):

What is the slope of the line described by the equation –4x + 6y = 12?

OpenStudy (anonymous):

first you have to get the intercepts of both x and y by plugging in 0 for one at a time. –4x + 6(0) = 12 –4x = 12 12 / -4 = x intercept. –4(0) + 6y = 12 6y = 12 12 / 6 = y intercept. can you do this for me?

OpenStudy (anonymous):

@marie123 ?

OpenStudy (anonymous):

um couldn't you just put this in slope intercept form?

OpenStudy (anonymous):

answer choices???

OpenStudy (anonymous):

if i'm right.... wouldn't you just get y by itself by adding x. and then dividing by 6???

OpenStudy (anonymous):

i'm not sure what you mean ^^ the easiest way for me is to convert then graph. i really don't understand any other method.